# Double-coset zeta functions for groups acting on trees

### v1

@@ 3.4. Local action diagrams and their associated universal groups
Following [17, Definition 3.1], a local action diagram is a triple
@@
    a connected graph $\Gamma$;
@@
    a family of non-empty pairwise disjoint sets $(X_{a})_{a\in E\Gamma}$. For every $c\in V\Gamma$, set $X_{c}=\bigsqcup_{a\in o^{-1}(c)}X_{a}$ and $X=\bigsqcup_{a\in E\Gamma}X_{a}$;
@@
If there is no ambiguity, we write $\Delta=(\Gamma,(X_{a}),(G(c)))$ in place of $\Delta=(\Gamma,(X_{a})_{a\in E\Gamma},(G(c))_{c\in V\Gamma})$. Moreover, given $u\in\Gamma$, set $X_{U}:=X_{u}$ if $u\in V\Gamma$ and $X_{U}:=X_{u}\sqcup X_{\bar{u}}$ if $u\in E\Gamma$.
Local action diagrams can be constructed from a group action on a tree $(G,T)$ as follows: Let $\pi$ be the quotient map on $(G,T)$ and choose a set of representatives $V^{*}$ of the $G$-orbits on $VT$. Following [17, Definition 3.6], the local action diagram associated to $(G,T)$ and $V^{*}$ is defined as follows:
@@
    $\Gamma=G\backslash T$ is the quotient graph of $(G,T)$;
@@
Note that the standard edge weight on the quotient graph $\Gamma$ is given by $\omega(a)=|X_{a}|$, for every $a\in E\Gamma$. Up to isomorphism of local action diagrams (cf. [17, Definition 3.2]), every group action on a tree $(G,T)$ has a unique associated local action diagram (cf. [17, Lemma 3.7]). Thus we refer to the local action diagram associated to $(G,T)$. Moreover, one of the key results in [17] (recalled in Theorem 3.8) shows that every local action diagram arises as the local action diagram associated to a group action on a tree.
@@
    Given a prime $p$, denote by $\mathbb{Q}_{p}$ the field of $p$-adic numbers, by $\mathbb{Z}_{p}$ the ring of $p$-adic integers, and by $\mathbb{F}_{p}\simeq\mathbb{Z}_{p}/{p\mathbb{Z}_{p}}$ the field of size $p$. Here below, we report and rephrase some results from [20, §II.1.4]. The action of $G=\mathrm{SL}_{2}(\mathbb{Q}_{p})$ on its Bruhat–Tits tree $T$ – which is a $(p+1)$-regular tree – has a $1$-segment as quotient graph $\Gamma$. Set $E\Gamma=\{a,\bar{a}\}$. Moreover, there is $e\in ET$ with $G\cdot e=a$ satisfying $G_{o(e)}=\mathrm{SL}_{2}(\mathbb{Z}_{p})\simeq G_{t(e)}$. Both the $\mathrm{SL}_{2}(\mathbb{Z}_{p})$-action on $\{f\in ET\mid o(f)=o(e)\}$ and the $G_{t(e)}$-action on $\{f\in ET\mid o(f)=o(\bar{e})\}$ are permutational isomorphic to the faithful action of $\mathrm{PSL}_{2}(\mathbb{F}_{p})$ on the projective line $\mathbb{P}^{1}(\mathbb{F}_{p})$. Hence, the local action diagram $\Delta=(\Gamma,(X_{b})_{b\in E\Gamma},(G(c))_{c\in V\Gamma})$ associated to $(G,T)$ is given by setting $X_{a}=X_{\bar{a}}=\mathbb{P}^{1}(\mathbb{F}_{p})$ and $G(o(a))=G(t(a))=\mathrm{PSL}_{2}(\mathbb{F}_{p})$.
We now expand the vocabulary of local action diagrams, introducing tools of key importance for the discussion.
@@
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram. For $n\geq 1$, an $n$-path in $\Delta$ is a sequence $\xi=(x_{1},\ldots,x_{n})$ obtained by starting with a path $\mathfrak{p}_{\xi}=(a_{1},\ldots,a_{n})$ in $\Gamma$ and selecting, for each $1\leq i\leq n$, an element $x_{i}\in X_{a_{i}}$. One says that $\xi$ starts at $x_{1}$, ends with $x_{n}$, has length $n$ (written $\ell(\xi)=n$), and $\mathfrak{p}_{\xi}$ is called the underlying path of $\xi$ in $\Gamma$. The 00-path $O_{c}$ in $\Delta$ at $c\in V\Gamma$ is the empty sequence of elements in $X$ with the trivial path at $c$ as underlying path in $\Gamma$. The path $O_{c}$ has length zero. A path in $\Delta$ is an $n$-path for some $n\geq 0$. Given paths $\xi=(x_{1},\ldots,x_{m})$ and $\eta=(y_{1},\ldots,y_{n})$ in $\Delta$ with $\mathfrak{p}_{\xi}=(a_{1},\ldots,a_{m})$ and $\mathfrak{p}_{\eta}=(b_{1},\ldots,b_{n})$, the product $\xi\cdot\eta$ is defined only if $t(a_{m})=o(b_{1})$ and it is the path $(x_{1},\ldots,x_{m},y_{1},\ldots,y_{n})$ in $\Delta$ with underlying path $\mathfrak{p}_{\xi}\cdot\mathfrak{p}_{\eta}$. Put also $\xi\cdot O_{t(a_{m})}=\xi$ and $O_{o(b_{1})}\cdot\eta=\eta$. Given a path $\xi$ and a non-empty set of paths $\mathscr{E}$ in $\Delta$ such that the product $\xi\cdot\eta$ is defined for every $\eta\in\mathscr{E}$, set $\xi\cdot\mathscr{E}=\{\xi\cdot\eta\mid\eta\in\mathscr{E}\}$. If $\xi=(x)$ has length $1$, we write $x\cdot\mathscr{E}$ in place of $(x)\cdot\mathscr{E}$.
A map $\iota\colon X\to X$ is said to be an inversion in $\Delta$ if $\iota(X_{a})\subseteq X_{\bar{a}}$ for every $a\in E\Gamma$. A path $\xi$ in $\Delta$ is reduced in $(\Delta,\iota)$ if it either $\ell(\xi)=0$ or $\xi=(x_{1},\dots,x_{n})$, for some $n\geq 1$, and $x_{i+1}\neq\iota(x_{i})$ for every $11\leq i\leq n-$ 1. Note that, even if $\xi$ is reduced, the underlying path $\mathfrak{p}_{\xi}$ needs not to be reduced. Denote by $\mathcal{P}_{(\Delta,\iota)}$ the set of all reduced paths in $(\Delta,\iota)$. For non-empty subsets $X_{1},X_{2}\subseteq X$, let also $\mathcal{P}_{(\Delta,\iota)}(X_{1}\to X_{2})$ be the collection of all reduced paths in $(\Delta,\iota)$ starting at some $x_{1}\in X_{1}$ and ending at some $x_{2}\in X_{2}$.
@@
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram. Following [17, Definition 3.4], a $\Delta$-tree $(T,\pi,\mathcal{L})$ consists of a tree $T$, a graph epimorphism $\pi\colon T\to\Gamma$ and a map $\mathcal{L}\colon ET\to X$ which restricts, for all $v\in VT$ and $a\in E\Gamma$ with $o(a)=\pi(v)$, to a bijection
@@
Note that the definition of a $\Delta$-tree is independent from $(G(c))_{c\in V\Gamma}$.
According to [6, §3.5], the pair $(T,\pi)$ is the universal cover of the edge-weighted graph $(\Gamma,\omega)$, where $\omega(a)=|X_{a}|$ for every $a\in E\Gamma$. Therefore, for every two $\Delta$-trees $(T,\pi,\mathcal{L})$ and $(T^{\prime},\pi^{\prime},\mathcal{L}^{\prime})$ there is a graph isomorphism $\varphi\colon T\to T^{\prime}$ such that $\pi=\pi^{\prime}\circ\varphi$ (cf. [17, Lemma 3.5]). Moreover, for every $\Delta$-tree we may define $\mathrm{Aut}_{\pi}(T)$ as in Example 1.1.
@@ 3.4.1. The standard $\Delta$-tree associated to $\iota$ and $c_{0}$
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram. Following the proof of [17, Lemma 3.5], we recall the construction of an explicit family of $\Delta$-trees which plays an important role in the next discussion.
@@
The vertex $v_{0}=O_{c_{0}}$ is called the root of $T$. The edges of $T$ are the pairs $(v,w)$ and $(w,v)$ of reduced paths in $\Delta$ of the form $v=(x_{1},\dots,x_{n})$ and $w=(x_{1},\dots,x_{n},x_{n+1})$, for some $n\geq 0$. Every edge $(v,w)$ of $T$ as before is said to be a positive edge. Denote by $ET^{+}$ the set of all positive edges of $T$. The origin, the terminus, and the inversion maps of $T$ are given by $o(v,w)=v$, $t(v,w)=w$ and $\overline{(v,w)}=(w,v)$, for every $(v,w)\in ET$. For every $(v,w)\in ET^{+}$ with $w=(x_{1},\dots,x_{n+1})$, set $\mathcal{L}(v,w)=x_{n+1}$ and $\mathcal{L}(w,v)=\iota(x_{n+1})$.
@@
By Remark 3.4.1, a path $\mathfrak{p}=(e_{1},\ldots,e_{n})$ in $T$ with $e_{1},\ldots,e_{n}\in ET^{+}$ is reduced, and only if, $\mathcal{L}(\mathfrak{p})$ is reduced in $(\Delta,\iota)$. Moreover, by (3.4), for every $v=(x_{1},\ldots,x_{n})\in VT$ there is a unique reduced path from $v_{0}$ to $v$ in $T$, namely $\mathcal{O}_{v_{0}}$ if $v=v_{0}$, and $(v_{0},e_{1},v_{1},\ldots,e_{n},v_{n}=v)$ with $v_{i}=(x_{1},\ldots,x_{i})$ for every $1\leq i\leq n$ otherwise. In particular, $T$ is connected.
@@
By (3.5), every edge lying in a geodesic from $v_{0}$ in $T$ belongs $ET^{+}$. Indeed, if $(e_{1},\ldots,e_{n})\in\mathrm{Geod}_{T}(v_{0}\to T)$ and $\mathcal{L}(e_{1},\ldots,e_{n})=(x_{1},\ldots,x_{n})$, then $e_{1}=(O_{c_{0}},x_{1})$ and $e_{i}=\big{(}(x_{1},\ldots,x_{i-1}),(x_{1},\ldots,x_{i})\big{)}$ for every $2\leq i\leq n$.
Moreover, if $e_{1}\in ET^{+}$, then for every $e_{2}\in o^{-1}(t(e_{1}))\setminus\{\bar{e}_{1}\}$ the path $[v_{0},e_{1}]\cdot e_{2}$ is a geodesic from $v_{0}$ and thus $e_{2}\in ET^{+}$.
We define the graph epimorphism $\pi\colon T\to\Gamma$ by putting $\pi(O_{c_{0}})=c_{0}$ and, provided $v=(x_{1},\ldots,x_{n})\in VT$ ($n\geq 1$) has underlying path $(a_{1},\ldots,a_{n})$ in $\Gamma$, by $\pi(v)=t(a_{n})$. The triple $(T,\pi,\mathcal{L})$ is a $\Delta$-tree that we call the standard $\Delta$-tree associated to $\iota$ and $c_{0}$.
@@
Write $e=(v,w)$, where $w=(x_{1},\ldots,x_{m},x)\in\mathcal{P}_{(\Delta,\iota)}(X_{c_{0}}\to X)$. Note that $\mathcal{L}([v_{0},v])=(x_{1},\ldots,x_{m})$. The bijection $\mathcal{L}^{v_{0}}$ in (3.6) restricts to a $1$-to-$1$ map
$[v_{0},v]\cdot\mathrm{Geod}_{T}(e\to T)\longrightarrow(x_{1},\ldots,x_{m})\cdot\mathcal{P}_{(\Delta,\iota)}(x\to X).$
@@
If $v=v_{0}$, then $o^{-1}(v)\subseteq ET^{+}$. From (3.9), (3.10) and the first part of the statement we conclude that $\mathcal{L}^{\bar{e}}$ is bijective. ∎
@@
Although it is not necessary for the discussion, for every $e=(v,w)\in ET^{+}$ one may restrict the map $\mathcal{L}$ in (3.5) to a bijection $\mathcal{L}^{\bar{e}}$ from $\mathrm{Geod}_{T}(\bar{e}\to T)$ to a suitable set of paths in $(\Delta,\iota)$. One may proceed inductively on $\ell([v_{0},v])=l\geq 0$. The case $l=0$ is done by Lemma 3.5. If $l\geq 1$, one assumes the claim true for $l-1$ and observes that $o^{-1}(v)\setminus\{e\}$ has exactly one edge which does not belong to $ET^{+}$: it is the edge whose reverse $f_{0}$ is the last edge of $[v_{0},v]$. By Lemma 3.5, for all $f\in o^{-1}(v)\setminus\{e,\bar{f}_{0}\}$, the map $\mathcal{L}^{f}$ as in (3.7) is bijective. Moreover, $\ell([v_{0},o(f_{0})])=\ell([v_{0},v])-1$ and by induction one has $\mathcal{L}^{\bar{f}_{0}}$ is bijective. Using the decomposition in (3.9), one determines the image of $\mathrm{Geod}_{T}(\bar{e}\to T)$ via $\mathcal{L}$ and thus the bijection $\mathcal{L}^{\bar{e}}$.

### v2

@@ 3.4. Local action diagrams and their associated universal groups
Following [24, Definition 3.1], a local action diagram is a triple
@@
    a connected graph $\Gamma$ in the sense [^1] I.e., $\Gamma$ is a connected graph in the sense of J-P. Serre, except that a loop may or may not be equal to its own reverse. of [24, §2.1];
@@
    a family of non-empty pairwise disjoint sets $(X_{a})_{a\in E\Gamma}$. For every $c\in V\Gamma$, set $X_{c}:=\bigsqcup_{a\in o^{-1}(c)}X_{a}$ and $X:=\bigsqcup_{a\in E\Gamma}X_{a}$;
@@
A local action diagram $\Delta=(\Gamma,(X_{a}),(G(c)))$ is trivial if $\Gamma$ is a $1$-point graph (i.e., a graph $\Gamma$ with $|V\Gamma|=1$ and $E\Gamma=\emptyset$).
In Theorem 3.8 we will recall a correspondence – established by C. Reid and S. Smith in [24] – between local action diagrams and (P)-closed actions on trees. In particular, trivial local action diagrams correspond to (P)-closed actions on $1$-point trees, and local action diagrams whose underlying graph has a $1$-loop equal to its own inverse correspond to (P)-closed actions on trees with edge inversions. Hence, having assumed Convention 3.0.1 for group actions on trees, we assume the following henceforth.
Convention 3.1.1.
In this paper, all local action diagrams are assumed to be non-trivial and their underlying graph is assumed to have no 1-loops equal to their own inverses (that is, it is graph as defined in Section 2.2).
@@
If there is no ambiguity, we write $\Delta=(\Gamma,(X_{a}),(G(c)))$ in place of $\Delta=(\Gamma,(X_{a})_{a\in E\Gamma},(G(c))_{c\in V\Gamma})$. Moreover, given $u\in\Gamma$, set $X_{U}:=X_{u}$ if $u\in V\Gamma$ and $X_{U}:=X_{u}\sqcup X_{\bar{u}}$ if $u\in E\Gamma$ (cf. Notation 2.1).
Local action diagrams can be constructed from a group action on a tree $(G,T)$ without edge inversions as follows: Let $\pi$ be the quotient map on $(G,T)$ and choose a set of representatives $V^{*}$ of the $G$-orbits on $VT$. Following [24, Definition 3.6], the local action diagram associated to $(G,T)$ and $V^{*}$ is defined as follows:
@@
    $\T\Gamma=G\backslash$ T is the quotient graph of $(G,T)$;
@@
Note that the standard edge weight on the quotient graph $\Gamma$ is given by $\omega(a)=|X_{a}|$, for every $a\in E\Gamma$. Up to isomorphism of local action diagrams (cf. [24, Definition 3.2]), every group action on a tree $(G,T)$ has a unique associated local action diagram (cf. [24, Lemma 3.7]). Thus, we refer to the local action diagram associated to $(G,T)$.
Note that two actions might have the same local action diagram (cf. Example 3.3(ii)). One of the key results in [24] (recalled in Theorem 3.8) shows that, up to natural notions of isomorphisms, there is a bijective correspondence between (P)-closed actions on trees and local action diagrams.
@@
    Given a prime $p$, denote by $\mathbb{Q}_{p}$ the field of $p$-adic numbers, by $\mathbb{Z}_{p}$ the ring of $p$-adic integers, and by $\mathbb{F}_{p}\simeq\mathbb{Z}_{p}/{p\mathbb{Z}_{p}}$ the field of size $p$. Here below, we report and rephrase some results from [27, §II.1.4].
    The action of $G=\mathrm{SL}_{2}(\mathbb{Q}_{p})$ on its Bruhat–Tits tree $T$, which is a $(p+1)$-regular tree, has a $1$-segment as quotient graph $\Gamma$. Set $E\Gamma=\{a,\bar{a}\}$. Moreover, there is $e\in ET$ with $G\cdot e=a$ satisfying $G_{o(e)}=\mathrm{SL}_{2}(\mathbb{Z}_{p})\simeq G_{t(e)}$. Both the $\mathrm{SL}_{2}(\mathbb{Z}_{p})$-action on $\{f\in ET\mid o(f)=o(e)\}$ and the $G_{t(e)}$-action on $\{f\in ET\mid o(f)=o(\bar{e})\}$ are permutational isomorphic to the faithful action of $\mathrm{PSL}_{2}(\mathbb{F}_{p})$ on the projective line $\mathbb{P}^{1}(\mathbb{F}_{p})$. Hence, the local action diagram $\Delta=(\Gamma,(X_{b})_{b\in E\Gamma},(G(c))_{c\in V\Gamma})$ associated to $(G,T)$ is given by setting $X_{a}=X_{\bar{a}}=\mathbb{P}^{1}(\mathbb{F}_{p})$ and $G(o(a))=G(t(a))=\mathrm{PSL}_{2}(\mathbb{F}_{p})$.
    The same local action diagram can be obtained by taking $G$ as the projective special linear group $\mathrm{PSL}_{2}(\mathbb{Q}_{p})$, and considering its induced action on $T$. Since the kernel of the $\mathrm{SL}_{2}(\mathbb{Q}_{p})$-action on $T$ is the centre of $\mathrm{SL}_{2}(\mathbb{Q}_{p})$, the induced $\mathrm{PSL}_{2}(\mathbb{Q}_{p})$-action on $T$ is faithful.
We now expand the vocabulary of local action diagrams, introducing some key tools for the discussion.
@@
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram as in Convention 3.1.1. For $n\geq 1$, an $n$-path in $\Delta$ is a sequence $\xi=(x_{1},\ldots,x_{n})$ obtained by starting with a path $\mathfrak{p}_{\xi}=(a_{1},\ldots,a_{n})$ in $\Gamma$ and selecting, for each $1\leq i\leq n$, an element $x_{i}\in X_{a_{i}}$. One says that $\xi$ starts at $x_{1}$, ends with $x_{n}$, has length $n$ (written $\ell(\xi)=n$), and $\mathfrak{p}_{\xi}$ is called the underlying path of $\xi$ in $\Gamma$. The 0-path $O_{c}$ in $\Delta$ at $c\in V\Gamma$ is the empty sequence of elements in $X$ with the trivial path at $c$ as underlying path in $\Gamma$. The path $O_{c}$ has length zero. A path in $\Delta$ is an $n$-path for some $n\geq 0$. Given paths $\xi=(x_{1},\ldots,x_{m})$ and $\eta=(y_{1},\ldots,y_{n})$ in $\Delta$ with $\mathfrak{p}_{\xi}=(a_{1},\ldots,a_{m})$ and $\mathfrak{p}_{\eta}=(b_{1},\ldots,b_{n})$, the product $\xi\cdot\eta$ is defined only if $t(a_{m})=o(b_{1})$ and it is the path $(x_{1},\ldots,x_{m},y_{1},\ldots,y_{n})$ in $\Delta$ with underlying path $\mathfrak{p}_{\xi}\cdot\mathfrak{p}_{\eta}$. Put also $\xi\cdot O_{t(a_{m})}=\xi$ and $O_{o(b_{1})}\cdot\eta=\eta$. Given a path $\xi$ and a non-empty set of paths $\mathscr{E}$ in $\Delta$ such that the product $\xi\cdot\eta$ is defined for every $\eta\in\mathscr{E}$, let $\xi\cdot\mathscr{E}=\{\xi\cdot\eta\mid\eta\in\mathscr{E}\}$. If $\xi=(x)$ has length $1$, we write $x\cdot\mathscr{E}$ in place of $(x)\cdot\mathscr{E}$.
A map $\iota\colon X\to X$ is said to be an inversion in $\Delta$ if $\iota(X_{a})\subseteq X_{\bar{a}}$ for every $a\in E\Gamma$. A path $\xi$ in $\Delta$ is reduced in $(\Delta,\iota)$ if either $\ell(\xi)=0$ or $\xi=(x_{1},\dots,x_{n})$, for some $n\geq 1$, and $x_{i+1}\neq\iota(x_{i})$ for every $1\leq i\leq n-1$. Even if $\xi$ is reduced, the underlying path $\mathfrak{p}_{\xi}$ needs not to be reduced. Denote by $\mathcal{P}_{(\Delta,\iota)}$ the set of all reduced paths in $(\Delta,\iota)$. For non-empty subsets $X_{1},X_{2}\subseteq X$, let also $\mathcal{P}_{(\Delta,\iota)}(X_{1}\to X_{2})$ be the collection of all reduced paths in $(\Delta,\iota)$ starting at some $x_{1}\in X_{1}$ and ending at some $x_{2}\in X_{2}$.
@@
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram. Following [24, Definition 3.4], a $\Delta$-tree $(T,\pi,\mathcal{L})$ consists of a tree $T$, a graph epimorphism $\pi\colon T\to\Gamma$ and a map $\mathcal{L}\colon ET\to X$ which restricts, for all $v\in VT$ and $a\in E\Gamma$ with $o(a)=\pi(v)$, to a bijection
@@
Note that the definition of a $\Delta$-tree is independent of $(G(c))_{c\in V\Gamma}$.
According to [8, §3.5], the pair $(T,\pi)$ is the universal cover of the edge-weighted graph $(\Gamma,\omega)$, where $\omega(a)=|X_{a}|$ for every $a\in E\Gamma$ (here $\omega(a)$ is regarded as a cardinal). Therefore, for every two $\Delta$-trees $(T,\pi,\mathcal{L})$ and $(T^{\prime},\pi^{\prime},\mathcal{L}^{\prime})$ there is a graph isomorphism $\varphi\colon T\to T^{\prime}$ such that $\pi=\pi^{\prime}\circ\varphi$ (cf. [24, Lemma 3.5]). Moreover, for every $\Delta$-tree we may define $\mathrm{Aut}_{\pi}(T)$ as in Example 1.1.
@@ 3.4.1. The standard $\Delta$-tree associated to $\iota$ and $c_{0}$
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram. Following the proof of [24, Lemma 3.5], we recall the construction of an explicit family of $\Delta$-trees which plays an important role in the next discussion.
@@
The vertex $v_{0}=O_{c_{0}}$ is called the root of $T$. The edges of $T$ are the pairs $(v,w)$ and $(w,v)$ of reduced paths in $\Delta$ of the form $v=(x_{1},\dots,x_{n})$ and $w=(x_{1},\dots,x_{n},x_{n+1})$, for some $n\geq 0$. Every edge $(v,w)$ of $T$ as before is said to be a positive edge. Denote by $ET^{+}$ the set of all positive edges of $T$. The origin, terminus, and inversion maps of $T$ are given by $o(v,w)=v$, $t(v,w)=w$ and $\overline{(v,w)}=(w,v)$, for every $(v,w)\in ET$. For every $(v,w)\in ET^{+}$ with $w=(x_{1},\dots,x_{n+1})$, set $\mathcal{L}(v,w)=x_{n+1}$ and $\mathcal{L}(w,v)=\iota(x_{n+1})$.
@@
By Remark 3.4.1, a path $\mathfrak{p}=(e_{1},\ldots,e_{n})$ in $T$ with $e_{1},\ldots,e_{n}\in ET^{+}$ is reduced if and only if $\mathcal{L}(\mathfrak{p})$ is reduced in $(\Delta,\iota)$. Moreover, by (3.4), for every $v=(x_{1},\ldots,x_{n})\in VT$ there is a unique reduced path from $v_{0}$ to $v$ in $T$, namely $\mathcal{O}_{v_{0}}$ if $v=v_{0}$, and $(v_{0},e_{1},v_{1},\ldots,e_{n},v_{n}=v)$ with $v_{i}=(x_{1},\ldots,x_{i})$ for every $1\leq i\leq n$ otherwise. In particular, $T$ is connected.
@@
By (3.5), every edge lying in a geodesic from $v_{0}$ in $T$ belongs to $ET^{+}$. Indeed, if $(e_{1},\ldots,e_{n})\in\mathrm{Geod}_{T}(v_{0}\to T)$ and $\mathcal{L}(e_{1},\ldots,e_{n})=(x_{1},\ldots,x_{n})$, then $e_{1}=(O_{c_{0}},x_{1})$ and $e_{i}=\big((x_{1},\ldots,x_{i-1}),(x_{1},\ldots,x_{i})\big)$ for every $2\leq i\leq n$.
Moreover, for $e\in ET^{+}$ and for every $f\in o^{-1}(t(e))\setminus\{\bar{e}\}$, the path $v_{0},e]\cdot f$ is a geodesic from $v_{0}$ and thus $f\in ET^{+}$.
We define the graph epimorphism $\pi\colon T\to\Gamma$ by putting $\pi(O_{c_{0}})=c_{0}$ and, for $v=(x_{1},\ldots,x_{n})\in VT$ ($n\geq 1$) having underlying path $(a_{1},\ldots,a_{n})$ in $\Gamma$, by $\pi(v)=t(a_{n})$. The triple $(T,\pi,\mathcal{L})$ is a $\Delta$-tree that we call the standard $\Delta$-tree associated to $\iota$ and $c_{0}$.
@@
Write $e=(v,w)$, where $w=(x_{1},\ldots,x_{m},x)\in\mathcal{P}_{(\Delta,\iota)}(X_{c_{0}}\to X)$. Note that $\mathcal{L}([v_{0},v])=(x_{1},\ldots,x_{m})$. The bijection $\mathcal{L}^{v_{0}}$ in (3.6) restricts to a bijective map
$v_{0},v]\cdot\mathrm{Geod}_{T}(e\to T)\longrightarrow(x_{1},\ldots,x_{m})\cdot\mathcal{P}_{(\Delta,\iota)}(x\to X).$
@@
If $v=v_{0}$, then $o^{-1}(v)\subseteq ET^{+}$. From (3.9), (3.10) and the first part of the statement, we conclude that $\mathcal{L}^{\bar{e}}$ is bijective. ∎
@@
For every $e=(v,w)\in ET^{+}$, one may restrict the map $\mathcal{L}$ in (3.5) to a bijection $\mathcal{L}^{\bar{e}}$ from $\mathrm{Geod}_{T}(\bar{e}\to T)$ to a suitable set of paths in $(\Delta,\iota)$. One may proceed inductively on $\ell([v_{0},v])=l\geq 0$. The case $l=0$ is done by Lemma 3.5. If $l\geq 1$, one assumes the claim true for l1l-1 and observes that $o^{-1}(v)\setminus\{e\}$ has exactly one edge which does not belong to $ET^{+}$: it is the edge whose reverse $f_{0}$ is the last edge of $v_{0},v]$. By Lemma 3.5, for all $f\in o^{-1}(v)\setminus\{e,\bar{f}_{0}\}$, the map $\mathcal{L}^{f}$ as in (3.7) is bijective. Moreover, $\ell([v_{0},o(f_{0})])=\ell([v_{0},v])-1$ and by induction $\mathcal{L}^{\bar{f}_{0}}$ is bijective. Using the decomposition in (3.9), one determines the image of $\mathrm{Geod}_{T}(\bar{e}\to T)$ via $\mathcal{L}$ and thus the bijection $\mathcal{L}^{\bar{e}}$.